OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x − ___

OpenStudy (saifoo.khan):

Since you know the x-intercepts.. You can find the whole equation..

OpenStudy (saifoo.khan):

x = -6 and x = 2 (x+6) = 0 and (x-2) =0 Now combine, (x+6)(x-2) = 0 Try using FOIL now..

OpenStudy (anonymous):

i am just not very good at working backwards

OpenStudy (anonymous):

what the heck is FOIL

OpenStudy (saifoo.khan):

Aluminium foil..

OpenStudy (anonymous):

ok i know about that i just hadn't heard the acronym before :)

OpenStudy (saifoo.khan):

Ah. Okay. The term is widely used in US though.

OpenStudy (anonymous):

i'm in the US and i still haven't heard of it :)

OpenStudy (saifoo.khan):

I see..